OpenStudy (anonymous):

How do I find the common algebraic expression. Example: Factor our the common algebraic expression raised to the lowest exponent Example: (x-3)^1/3 - (x-3) ^-2/3 = 0

OpenStudy (amistre64):

since -2/3 is the lowest; divide off (x-3)^2/3

OpenStudy (amistre64):

err, ^-2/3

OpenStudy (amistre64):

\[ (x-3)^{1/3} - (x-3)^{-2/3} = 0\] \[ \frac{(x-3)^{1/3}}{(x-3)^{-2/3}} -\frac{(x-3)^{-2/3}}{(x-3)^{-2/3}} = \frac{0}{(x-3)^{-2/3}}\] \[ (x-3)^{2/3} ((x-3) - 1) = 0\]

OpenStudy (amistre64):

from there I got no idea what its asking for :/